OpenStudy (anonymous):

simplify completely: 6x^4+9x^2+12x ---------------- 3x Answer: 2x^5+3x^3+4x^2???

OpenStudy (wasiqss):

6x^4/3x=2x^3 9x^2/3x=3x 12x/3x=4 well final answer is 2x^3+3x+4

OpenStudy (wasiqss):

yehh

OpenStudy (anonymous):

oh crap! yea!! thanks! (: